[Brahms-l] Shift report 20040428 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Wed Apr 28 2004 - 07:44:06 EDT
------------------------------------------------------------
Date: Wed Apr 28 07:44:06 2004, Shift: 00:00-08:00
Supervisor: EBJ, with: 
Log book#15 pages 115 - 115

Shift Summary:
------------------------------------------------------------
1. RHIC was down for most of the night.  We have beam now and are collecting data.

2. Finished off the FS 3 deg 1/sqrt(2) A data set 
   Finished off the MRS 90 deg 175 B data set

3. I didn't discover any new detector problems.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Apr 28 07:44:22 2004

This archive was generated by hypermail 2.1.8 : Wed Apr 28 2004 - 07:44:47 EDT